Operations Report – January 16th

Supplemental Operations Report 16Jan2022

Name of person filing report: Atila Meszaros

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Total run time 424.1 hours. Still being manually run at night. 76 more hours for the generator’s air filter and spark plug maintenance, parts have been already received. Generator’s oil filter maintenance was done today 100 hours after the last oil change. Oil filter was replaced and a new one was installed. Oil level is full.

ScienceDome Dual Split: Off

Solar—Nominal, charge controllers off at night. We turn the generator on before the SOC reaches 70%.

Solar— VDC Last daylight hours: (information obtained from

Average 51.35

Minimum 46.8

Maximum 59.3

Notes on solar: Nothing to report

Propane Reading, station tank – 25 %

Propane Reading, director tank— 30 %

Propane Reading, intern tank— 33 %

Propane Reading, generator— 65 %

Ethanol Free Gasoline – 4 gallons

Water (Outpost tank) – 300 gallons

Perseverance rover used: No

Hours: 226.1

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Sojourner rover used: No

Hours: 160.8

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: All rovers batteries have been checked and every cell has been refilled. Curiosity’s low performance on Thursday was definitely due to its battery’s cells being almost empty. All of the rovers needed refilled, except Percy, which was refilled some weeks ago. However, the batteries’ connection showed acid on the surface and has been deeply cleaned with baking soda.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: Oil checked on both, Honda and 350.1 again. Oil levels nominal.

HabCar used and why, where? Yes, for water runs for both Greenhouse and Hab’s static tank. Oil, transmission and coolant levels are looking good.

CrewCar used and why, where? Yes, to transport the last crew to GJT. Still awaiting for the new crew’s PCR covid test to have clearance to arrive at MDRS.

Luna used and why, where? No

Business internet: Nothing to report

General notes and comments: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: New propane tank for the generator is installed. It is 500 gallons (400 worth of propane), so it is double the size of the one we had before. Blackburn will come early on the week to refill all our 4 propane tanks. Science Dome foundation has been secured, all the structure is surrounded with gravel taken from the vicinity of the observatory. Furnace and lower deck space has been measured and plotted for new furnace installation. Generator still turned on every evening around 8 PM at a 70-80% SOC. Awaiting for new toilet installation in the next few days. Tomorrow will take the CrewCar to Ryan’s Auto for service and oil change. Fertilization of our crops is due tomorrow. 1 helmet of the previous report has additional scratches on the right side. Seems that it has a bit of white paint on it. Two helmets of suits 7 and 8 present grey scratches above forehead level (probably from hitting the Rover top part). CO monitors arrived and has been placed in the RAM. All the food and supplies for the crew is already on campus, all the water tanks are full and we are looking forward to having them here!

Summary of Outpost operations: Intern Trailer back side had some uncovered holes that have been gorilla taped after cleaning the surface from old tape. A section under the trailer was detected to be open, probably causing the freezing of our water pipes during cold nights, has been partially secured with clamps, but drilling is needed to secure the piece. Two jacks have been secured under the Intern Trailer in what would be the living room area. It is definitely more stable. Both plants from the Science Dome have been moved to the Intern Trailer.

Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site